R-134a flows in a pipe at 30oC with a specific volume of 0.04434 m3/kg. The internal energy and enthalpy of R-134a are 250.8 kJ/kg and 273.0 kJ/kg, respectively. Determine the pressure of the refrigerant in MPa. in 2 decimal places


1
Expert's answer
2021-08-16T08:32:48-0400

"V=0.04434 \;m^3/kg \\
U = 250.8 \;kJ/kg \\
h= 273.0 \;kJ/kg \\
h=U+pV \\
273.0 = 250.8 + P \times 0.04434 \\
P = 500.67 \;kPa = 0.50 \;MPa"